# Stability of the expanding region of Kerr de Sitter spacetimes

### v1

@@ 1 Introduction
where ${\bm{\mathrm{Ric}}}[{\bm{g}}]$ is the Ricci curvature of the spacetime metric ${\bm{g}}$ on a $3+1$-dimensional Lorentzian manifold $(\mathcal{M},{\bm{g}})$. Since the sum of the sectional curvatures $\bm{K}_{i}$ is negative, [^1] Here $(e_{0},e_{1},e_{2},e_{3})$ is an orthonormal frame, $e_{0}$ is time-like, and $\bm{K}_{i}=\bm{R}(e_{0},e_{i},e_{0},e_{i})$ are the sectional curvatures associated to the planes spanned by $e_{0}$ and $e_{i}$.
${\bm{\mathrm{Ric}}}[{\bm{g}}](e_{0},e_{0})=\sum_{i=1}^{3}\,\bm{K}_{i}=\Lambda\>{\bm{g}}(e_{0},e_{0})<0$
solutions to (1.1) may exhibit expansion in all directions. The simplest example of a spacetime with this property is de Sitter space, which models a spatially closed expanding universe with topology $\mathbb{S}^{3}\times\mathbb{R}$. The exact solutions to (1.1) that motivate this paper are the Kerr de Sitter spacetimes $(\mathcal{M},{\bm{g}}_{\mathcal{K}_{a,m}})$. In addition to black hole interior and exterior regions, they contain a spatially open expanding, or cosmological region, with topology ×$\mathbb{R}\times\mathbb{S}^{2}\times\mathbb{R}$ [^2] Carter gives an excellent discussion of the maximal extension of Kerr de Sitter in [Car09]. For an introduction to the global geometry of the cosmological region, specifically in the context of the Cauchy problem see [Sch15, Sch22].
@@
Small perturbations on ×$\mathbb{R}\times\mathbb{S}^{2}$ which converge exponentially fast at both ends to nearby members of the Kerr de Sitter family lead to future geodesically complete solutions to (1.1) which display asymptotically de Sitter-like degrees of freedom.
@@
Recall the Penrose diagram of the Kerr de Sitter metric ${\bm{g}}_{\mathcal{K}_{a,m}}$ in Fig. 1, for small angular momentum $a^{2}\ll m^{2}\ll\Lambda^{-4}$. The cosmological region $\mathcal{R}$ lies to the future of the black hole exteriors $\mathcal{S}_{1}$, and $\mathcal{S}_{2}$, and is separated from these by the cosmological horizons $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$. The conformal boundary at infinity is denoted by $\mathcal{I}^{+}$, and the black hole regions $\mathcal{B}_{1}$, and $\mathcal{B}_{2}$ are in the complement of its past.
@@
$\Phi_{\mathcal{K}_{a,m}}^{2}=1+\mathcal{O}(e^{-2Hs})\,,\qquad(g_{s,\mathcal{K}_{a,m}})_{ij}=\mathcal{O}(e^{2Hs})\,;$ (1.6)
[…]
@@
${\bm{g}}-{\bm{g}}_{\mathcal{K}_{a_{i},m_{i}}}=\mathcal{O}(e^{-\alpha t})\quad\text{: in }\mathcal{S}_{i}\,,\qquad i=1,2\,,$ (1.7)
for some parameters $(a_{1},m_{1})$ and $(a_{2},m_{2})$ with $\sum_{i=1}^{2}|a_{i}|+|m_{i}-m|\ll 1$, and $\alpha>0$. In fact, the stability result (1.7) holds on a domain that goes beyond the event and cosmological horizons, uniformly in $r$ [HV18, Theorem 1.1]. Therefore the existence of a development is known on a domain (whose future boundary is indicated by the dash dotted lines in Figure 2) which contains a level set $\Sigma_{s_{0}}\simeq\mathbb{R}\times\mathbb{S}^{2}$ of $r$, with the property that the geometric data on $\Sigma_{s_{0}}$ converges along one end to that induced by ${\bm{g}}_{\mathcal{K}_{a_{1},m_{1}}}$, and that induced by ${\bm{g}}_{\mathcal{K}_{a_{2},m_{2}}}$ along the other. This is the initial data for the evolution problem we consider.
@@
with each leaf $\Sigma_{s}$ being diffeomorphic to a cylinder ×$\mathbb{R}\times\mathbb{S}^{2}$. We choose coordinates so that
${\bm{g}}=-\Phi^{2}\mathrm{d}s^{2}+(g_{s})_{ij}\mathrm{d}x^{i}\mathrm{d}x^{j}$ (1.9)
@@
Given the differentiable structure of $\mathcal{R}$, we can view ${\bm{g}}_{\mathcal{K}_{a,m}}$ as a family of metrics on $\mathcal{R}$. Moreover
@@
where $g$, $\widetilde{g}$, and $k$, $\widetilde{k}$ are the first and second fundamental forms of ${\bm{g}}$, and $\widetilde{{\bm{g}}}$ on $\Sigma_{s}$, respectively.
@@
$\mathcal{E}_{N}(s)=\|\widehat{g}\|^{2}_{H^{N}(\Sigma_{s},g)}+\|\widehat{g}^{-1}\|^{2}_{H^{N}(\Sigma_{s},g)}+e^{3Hs}\|\widehat{\Phi}\|^{2}_{H^{N}(\Sigma_{s},g)}+e^{2Hs}\big{(}\|\widehat{\Gamma}\|^{2}_{H^{N}(\Sigma_{s},g)}+\|\widehat{k}\|^{2}_{H^{N}(\Sigma_{s},g)}\big{)}$ (1.13)
@@
Moreover ×$\mathbb{R}\times\mathbb{S}^{2}$ is endowed with the standard metric on the cylinder,
@@
Theorem 1.
Let $(\mathcal{R},{\bm{g}})$ be a solution to (1.1) in parabolic gauge, namely expressed in coordinates (1.9) relative to a time function $s:\mathcal{R}\to(0,\infty)$ whose level sets have topology ×$\mathbb{R}\times\mathbb{S}^{2}$ and satisfy the geometric gauge condition (1.12).
Suppose for some $\mathring{\varepsilon}>0$ and $s_{0}>0$, the initial data $(g_{0},k_{0})$ on $\Sigma_{s_{0}}$ is sufficiently close to the data induced by a Kerr de Sitter metric ${\bm{g}}_{\mathcal{K}_{a,m}}$ expressed in this gauge, with parameters $(a_{1},m_{1})$ at one end, and $(a_{2},m_{2})$ at the other end, in the sense that for some $N\geq 4$, with the energy defined in (1.13):
@@
    Then, for $\mathring{\varepsilon}>0$ sufficiently small, the solution is global,
    $\mathcal{R}=\bigcup_{s=s_{0}}^{\infty}\Sigma_{s}\,,\qquad\text{and}\quad\mathcal{E}_{N}(s)\leq C\>\mathcal{E}_{N}(s_{0})\qquad(s\geq s_{0})\,.$ (1.16)
@@
    where $\widetilde{g}^{\infty}$, $g^{\infty},h,\widetilde{h}$ are metrics on ×$\mathbb{R}\times\mathbb{S}^{2}$, with $\widetilde{g}^{\infty}_{ij},\widetilde{h}_{ij}$ induced by the reference metric, satisfying
@@
Figure 3: Topology of the the level sets $\Sigma_{s}$ diffeomorphic to ×$\mathbb{R}\times\mathbb{S}^{2}$.
@@
The exponential decay assumption is not necessary: The theorem holds with $\alpha_{1}=\alpha_{2}=0$. We have included the exponential decay assumption to show that if the perturbation decays at the level of the initial data, then exponential decay is inherited along every $\Sigma_{s}$: This is true in particular for the asymptotics (1.18), (1.20). Exponential decay in Kerr de Sitter along the cosmological horizon is well-established for small angular momentum [^10] We remark that our theorem does not rely on a smallness assumption of $a_{i}$, $i=1,2$. If the black hole exteriors $\mathcal{S}_{i}$ were proven to be stable in the whole subextremal range, then our theorem provides an immediate extension to the cosmological region. in various settings: For the linear wave equation [BH08, Dya11b, Dya11a, Mav23], quasi-linear equations [Hin16, HV16, Hin17, Mav24], and – justifying our assumption here – for the Einstein equations [HV18, Fan22a, Fan22b].
@@
The fact that it is possible to prove global existence with a fixed reference metric $\widetilde{{\bm{g}}}$ means that on every time slice $\Sigma_{s}$, $s\geq s_{0}$, the solution tends to ${\bm{g}}_{\mathcal{K}_{a,m}}$ with the same parameters – $(a_{1},m_{1})$ towards $\iota^{+}_{1}$, and $(a_{2},m_{2})$ towards $\iota^{+}_{2}$, see Fig 2. This is markedly different from the proofs of the major black hole stability theorems [HV18, KS20, DHRT21, GKS22, KS23]: In the $\Lambda>0$ case, an iterative scheme for the linearised equations is implemented in [HV18] to make successive gauge corrections and to find the parameters $(a,m)$ of the final state; in the case $\Lambda=0$ the modulation techniques [KS22a, KS22b, She23] are used in [KS23] both to anchor the gauge, and to determine the parameters $(a,m)$ of the final state. Heuristically, the reason that the parameters of the Kerr de Sitter metric remain unchanged in the cosmological region is that they are only relevant for convergence along spacelike hypersurfaces — which at their endpoints are unaffected by the perturbation.
@@
The main asymptotic degrees of freedom are functional in nature: [^11] This is already the case for the wave equation [Vas10, Sch15], and gives rise to a scattering problem [Ber24]. the leading orders of the solution ${\bm{g}}$ in $s$ are not captured by a member ${\bm{g}}_{\mathcal{K}_{a,m}}$ of the Kerr de Sitter family, but given by a free function, as in (1.17).  [^12] Yet locally in the past of a given point on the conformal boundary the solution is asymptotically de Sitter. Cf. discussions of the cosmic no hair conjecture in [AR16, Sch22]; for a proof in spherical symmetry see [CNO19]. The significance of this effect for the theory of gravitational radiation has been suggested by Ashtekar et al [ABK16]. We refer in this context also to the gluing and scattering constructions [Hin21, Hin24, GVK17].
@@
A significant aspect of our theorem is that the topology of the spatial slices $\Sigma_{s}$ is ×$\mathbb{R}\times\mathbb{S}^{2}$ (and thus in the conformal diagram of Figure 2 the hypersurfaces $\Sigma_{s}$ extend to the end points $\iota_{1}^{+}$ and $\iota_{2}^{+}$, see Figure 3). Indeed for compact subsets $K\subset\Sigma_{s}$, with $s\geq s_{0}$ taken sufficiently large, a range of results in the literature imply that the domain of dependence $\mathcal{D}^{+}(K)$ is contained in a perturbation of de Sitter space: [^13] For a longer discussion of the spatially compact setting and its relation to earlier work on de Sitter see [Sch22, Section 1.6]. The de Sitter solution $(\mathbb{H},h)$ can be realized as a hyperboloid $\mathbb{H}$ in $\mathbb{R}^{4+1}$, with metric $h=m\rvert_{\mathbb{H}}$ induced by the ambient Minkowski metric $m$. See also [Vas10] or [Sch21] where the embedding $\mathbb{H}\subset\mathbb{R}^{4+1}$ is used, and its geometric properties are further discussed. In particular the conformal method applies [Fri86, MVK23], and more generally, Ringström proved geodesic completeness from spatially bounded data irrespective of the topology [Rin08].
@@
The gauge (1.12) is crucial for the global existence proof in this setting and in particular leads to a foliation that correctly identifies $\mathcal{I}^{+}$. A similar problem occurs in the study of stable spacelike singularities, where related ADM gauges have been used to synchronize the singularity [RS18, FRS23, FL23]; see also [AF20]. In the setting of expanding cosmologies, related gauges have been used in [LvEUW04, Fou22, FMO24].
@@
Since [HV18] provides the stability of the regions $\mathcal{S}_{i}$ on a domain that extends beyond both the cosmological horizon and the event horizon, the result of Hintz-Vasy can in principle also be combined with a theorem of Dafermos-Luk on the $C^{0}$-stability of the Cauchy horizon [DL17, Section 1.6]. [^14] For results in spherical symmetry with $\Lambda>0$ in this region see [CGaNS15a, CGaNS15b, CGaNS17]. The combination of all three theorems shows in particular that the Penrose diagram in Fig. 1 is dynamically stable.

### v2

@@ 1 Introduction
where ${\bm{\mathrm{Ric}}}[{\bm{g}}]$ is the Ricci curvature of the spacetime metric $\bm{g}$} on a $1+3$-dimensional Lorentzian manifold $(\mathcal{M},{\bm{g}})$. Since the sum of the sectional curvatures $\bm{K}_{i}$ is negative, [^1] Here $(e_{0},e_{1},e_{2},e_{3})$ is an orthonormal frame, $e_{0}$ is timelike, and $\bm{K}_{i}=\bm{R}(e_{0},e_{i},e_{0},e_{i})$ are the sectional curvatures associated to the planes spanned by $e_{0}$ and $e_{i}$.
${\bm{\mathrm{Ric}}}[{\bm{g}}](e_{0},e_{0})=\sum_{i=1}^{3}\,\bm{K}_{i}=\Lambda\>{\bm{g}}(e_{0},e_{0})<0\,$,
solutions to (1.1) may exhibit expansion in all directions. The simplest example of a spacetime with this property is de Sitter space, which models a spatially closed expanding universe with topology $\mathbb{R}\times\mathbb{S}^{3}$. The exact solutions to (1.1) that motivate this paper are the Kerr de Sitter spacetimes $(\mathcal{M},{\bm{g}}_{\mathcal{K}_{a,m}})$. In addition to black hole interior and exterior regions, they contain a spatially open expanding, or cosmological region, with topology $\mathbb{R}\times\mathbb{S}^{2}\times\mathbb{R}$. [^2] Carter gives an excellent discussion of the maximal extension of Kerr de Sitter in [Car09]. For an introduction to the global geometry of the cosmological region, specifically in the context of the Cauchy problem see [Sch15, Sch22].
@@
Small perturbations on $\mathbb{R}\times\mathbb{S}^{2}$ which converge exponentially fast at both ends to nearby members of the Kerr de Sitter family lead to future geodesically complete solutions to (1.1) which display asymptotically de Sitter-like degrees of freedom.
@@
Recall the Penrose diagram of the Kerr de Sitter metric ${\bm{g}}_{\mathcal{K}_{a,m}}$ in Fig. 1, for small angular momentum $a$, $a^{2}\ll m^{2}\ll\Lambda^{-1}$. The cosmological region $\mathcal{R}$ lies to the future of the black hole exteriors $\mathcal{S}_{1}$, and $\mathcal{S}_{2}$, and is separated from these by the cosmological horizons $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$. The conformal boundary at infinity is denoted by $\mathcal{I}^{+}$, and the black hole regions $\mathcal{B}_{1}$, and $\mathcal{B}_{2}$ are in the complement of its past.
@@
$\Phi_{\mathcal{K}_{a,m}}^{2}=1+\mathcal{O}(e^{-2Hs})\,,\hskip 18.49988pt(g_{s,\mathcal{K}_{a,m}})_{ij}=\mathcal{O}(e^{2Hs})\,;$ (1.6)
for explicit formulas see Section 3.1. Schwarzschild de Sitter is obtained by setting $a=0$. In fact, $g_{\mathcal{K}_{a,m}}$ is a metric on $\mathbb{R}\times\mathbb{S}^{2}$ and there are coordinates $(x^{1},x^{2},x^{3})=(t,x^{2},x^{3})$ such that $(g_{\mathcal{K}_{a,m}})_{ij}$ is independent of the $t$, and $(x^{2},x^{3})$ are spherical coordinates on $\mathbb{S}^{2}$. In other words, the metric $g_{\mathcal{K}_{a,m}}$ is translation invariant along the cylinder, and for $a=0$ takes the form: [^5] See [Sch15, Section 3, 5 & Appendix B] and [GH77, Sch15] for a detailed discussion of the geometry of the cosmological region, in these and comoving coordinates.
$g_{\mathcal{K}_{m,0}}=\Bigl(\frac{\Lambda r^{2}}{3}+\frac{2m}{r}-1\Bigr)\mathrm{d}t^{2}+r^{2}\mathrm{d}\sigma^{2}_{\mathbb{S}^{2}}$ (1.7)
@@
Theorem 1.1 (Stability of the stationary black hole exterior of slowly rotating Kerr de Sitter, Theorem 1.1 in [HV18]).
For smooth initial data $(g,k)$, close to the data $(g_{\mathcal{K}_{0,m}},k_{\mathcal{K}_{0,m}})$ induced by the Schwarzschild de Sitter metric ${\bm{g}}_{\mathcal{K}_{0,m}}$ on $\Sigma$ (cf. Figure 2) in a higher Sobolev norm, there exists a solution to (1.1) attaining the prescribed initial data on $\Sigma$, and parameters $(a_{1},m_{1})$ and $(a_{2},m_{2})$ with $\sum_{i=1}^{2}|a_{i}|+|m_{i}-m|\ll 1$, so that
${\bm{g}}-{\bm{g}}_{\mathcal{K}_{a_{i},m_{i}}}=\mathcal{O}(e^{-\alpha_{i}t})\quad\text{: in }\mathcal{S}_{i}\,,\qquad i=1,2\,$, (1.8)
for constants $\alpha_{i}>0$, $i=1,2$ independent of the initial data.
In fact, the stability result (1.8) holds on a domain that goes beyond the event and cosmological horizons, uniformly in $r$; see [HV18, Theorem 1.4] for precise statements, and Section 3.5 therein for a discussion of the extension beyond the horizon. Therefore the existence of a development is known on a domain (whose future boundary is indicated by the dash dotted lines in Figure 2) which contains a level set $\Sigma_{s_{0}}\simeq\mathbb{R}\times\mathbb{S}^{2}$ of $r$, with the property that the geometric data on $\Sigma_{s_{0}}$ converges along one end to that induced by ${\bm{g}}_{\mathcal{K}_{a_{1},m_{1}}}$, and that induced by ${\bm{g}}_{\mathcal{K}_{a_{2},m_{2}}}$ along the other. This is the initial data for the evolution problem we consider.
@@
with each leaf $\Sigma_{s}$ being diffeomorphic to a cylinder $\mathbb{R}\times\mathbb{S}^{2}$. We choose coordinates so that
${\bm{g}}=-\Phi^{2}\mathrm{d}s^{2}+(g_{s})_{ij}\mathrm{d}x^{i}\mathrm{d}x^{j}\,$, (1.10)
@@
In fact, in coordinates $(x^{1},x^{2},x^{3})=(t,x^{2},x^{3})$ on $\mathbb{R}\times\mathbb{S}^{2}$, the metric components of the solution $\bm{g}$} as well as of any member ${\bm{g}}_{\mathcal{K}_{a,m}}$ of the Kerr de Sitter family can be viewed as functions of $(r,t,x^{2},x^{3})$ in the fixed charts for the expanding region of Schwarzschild de Sitter. Then also
@@
where $g$, $\widetilde{g}$, and $k$, $\widetilde{k}$ are the first and second fundamental forms of $\bm{g}$}, and $\widetilde{{\bm{g}}}$ on $\Sigma_{s}$, respectively. In these coordinates, we then also obtain from (1.8) that for $s=s_{0}$, $\widehat{\Phi},\widehat{g},\widehat{k}\to 0$ as $t\to\pm\infty$; cf. Figure 3.
@@
[…]
@@
Moreover $\mathbb{R}\times\mathbb{S}^{2}$ is endowed with the standard metric on the cylinder, and in coordinates $(x^{1},x^{2},x^{3})=(t,\theta,\phi)$,
@@
Main theorem.
We are now ready to state the main result in this paper, which is a global existence result for the cosmological region, for initial data close to Kerr de Sitter. For local existence in parabolic gauge see Appendix A.
Theorem 1.2 (Stability of the expanding region of Kerr de Sitter).
Let $(\mathcal{R},{\bm{g}})$ be a local in time solution to (1.1) in parabolic gauge, expressed relative to a time function $s$ whose level sets satisfy the geometric gauge condition (1.13), and
$\mathcal{R}=\bigcup_{s=s_{0}}^{s_{1}}\Sigma_{s}\,,\qquad\Sigma_{s}\simeq\mathbb{R}\times\mathbb{S}^{2}\,$,
for some $s_{1}>s_{0}>0$, with $\bm{g}$} expressed in local coordinates as in (1.10).
For $i=1,2$, let $|a_{i}|\ll m_{i}\ll\Lambda^{-1/2}$, and $\alpha_{i}\geq 0$. Suppose for some $\mathring{\varepsilon}>0$, the initial data $(g_{0},k_{0})$ on $\Sigma_{s_{0}}$ is $\mathring{\varepsilon}$-close to the data induced by a Kerr de Sitter metric ${\bm{g}}_{\mathcal{K}_{a,m}}$ expressed in this gauge, with parameters $(a_{1},m_{1})$ at one end, and $(a_{2},m_{2})$ at the other end, in the sense that for some $N\geq 4$, with the energy defined in (1.14):
@@
    Then, for $|a_{1}-a_{2}|$, $|m_{1}-m_{2}|$, and $\mathring{\varepsilon}>0$ sufficiently small, the solution is global,
[…]
@@
    where $\widetilde{g}^{\infty}$, $g^{\infty},h,\widetilde{h}$ are metrics on $\mathbb{R}\times\mathbb{S}^{2}$, with $\widetilde{g}^{\infty}_{ij},\widetilde{h}_{ij}$ induced by the reference metric, satisfying
@@
Figure 3: Topology of the the level sets $\Sigma_{s}$ diffeomorphic to $\mathbb{R}\times\mathbb{S}^{2}$.
The stability of the expanding region of Kerr de Sitter established in Theorem 1.2 is a self-contained result for the Einstein equations (1.1), and independent of the stability of the stationary black hole exterior given by Theorem 1.1. Indeed for any smooth Cauchy data on $\Sigma_{s_{0}}$ sufficiently close to Kerr de Sitter satisfying (1.16), the solution is global to the future as described in Theorem 1.2. In particular, this result does not require the data to converge exponentially to Kerr de Sitter: Theorem 1.2 holds with $\alpha_{1}=\alpha_{2}=0$. Nor does it rely on a smallness assumption for the angular momenta $a_{i}$, i=1,2i=1,2.101010In particular, if the black hole exteriors $\mathcal{S}_{i}$ were proven to be stable in the whole subextremal range, then Theorem 1.2 provides an immediate extension to the cosmological region. However for the global stability problem for slowly rotating Kerr de Sitter spacetimes, Theorem 1.1 implies that the assumptions of Theorem 1.2 are satisfied for some $(a_{i},m_{i})$, $\alpha_{i}>0$, $i=1,2$, yielding a global nonlinear stability result, in the domain of dependence of a Cauchy hypersurface bridging two black hole exteriors:
Corollary 1.3.
For smooth initial data $(g,k)$ on $\Sigma$ as in Figure 2, close to the initial data induced by the Schwarzschild de Sitter metric as in Theorem 1.1, there exists a solution to (1.1) on $\mathcal{S}_{1}\cup\mathcal{R}\cup\mathcal{S}_{2}$, [^11] Here $\mathcal{S}_{i}$, $i=1,2$ are diffeomorphic to the domains to the future of $\Sigma$ in Schwarzschild de Sitter, containing a small extension of the stationary black hole exteriors, uniformly in $r$; [HV18, Section 3.5]. attaining the prescribed Cauchy data on $\Sigma$, which satisfies the asymptotic stability statement (1.8) in $\mathcal{S}_{i}$, $i=1,2$, and $(\mathcal{R},g)$ is global in the sense of Theorem 1.2 (I). Moreover, the solution satisfies the orbital stability (towards $\mathcal{I}^{+}$) and asymptotic stability (towards $\iota^{+}_{i}$, $i=1,2$) statements of Theorem 1.2 (II,III) with $\alpha_{i}>0$, $i=1,2$ as in (1.8).
@@
The estimates (1.19) and (1.21) show in particular that exponential decay established in (1.8) along $\mathcal{C}_{i}$, $i=1,2$, is inherited along every $\Sigma_{s}$ towards $\iota_{i}^{+}$, $i=1,2$. Exponential decay in Kerr de Sitter along the cosmological horizon is well-established for small angular momentum in various settings: For the linear wave equation [BH08, Dya11b, Dya11a, Mav23], quasi-linear equations [Hin16, HV16, Hin17, Mav24], and – justifying our assumption here – for the Einstein equations [HV18, Fan22a, Fan22b].
@@
The fact that it is possible to prove global existence with a fixed reference metric $\widetilde{{\bm{g}}}$ means that on every time slice $\Sigma_{s}$, $s\geq s_{0}$, the solution tends to ${\bm{g}}_{\mathcal{K}_{a,m}}$ with the same parameters – $(a_{1},m_{1})$ towards $\iota^{+}_{1}$, and $(a_{2},m_{2})$ towards $\iota^{+}_{2}$, see Fig 2. This is markedly different from the proofs of the major black hole stability theorems [HV18, DHR19, KS20, DHRT21, GKS24, KS23]: In the $\Lambda>0$ case, an iterative scheme for the linearised equations is implemented in [HV18] to make successive gauge corrections and to find the parameters $(a,m)$ of the final state; in the case $\Lambda=0$ the modulation techniques [KS22a, KS22b, She23] are used in [KS23] both to anchor the gauge, and to determine the parameters $(a,m)$ of the final state. Heuristically, the reason that the parameters of the Kerr de Sitter metric remain unchanged in the cosmological region is that they are only relevant for convergence along spacelike hypersurfaces — which at their endpoints are unaffected by the perturbation.
@@
The main asymptotic degrees of freedom are functional in nature: [^12] This is already the case for the wave equation [Vas10, Sch15], and gives rise to a scattering problem [Ber25]. The leading orders of the solution $\bm{g}$} in $s$ are not captured by a member ${\bm{g}}_{\mathcal{K}_{a,m}}$ of the Kerr de Sitter family, but given by a free function, as in (1.18).  [^13] Yet locally in the past of a given point on the conformal boundary the solution is asymptotically de Sitter. Cf. discussions of the cosmic no hair conjecture in [AR16, Sch22]; for a proof in spherical symmetry see [CNO19]. These functions are free up to constraints which take a particularly simple form on the conformal boundary [Ren04, Kro16]. The significance of this freedom for the theory of gravitational radiation has been suggested by Ashtekar et al [ABK16]. We remark, however, that the free functions $g_{ij}^{\infty}(x)$ in the expansion (1.18) capture only part of the asymptotic data set on the conformal boundary. The remaining free functions appear in higher order expansions of the metric, in particular, at order $\mathcal{O}(e^{-Hs})$ for $g_{ij}(s,x)$; (see also Remark 6.2). In the setting of FLRW spacetimes this is obtained in [Fou22, Theorem 1.3]. More recently, Hintz-Vasy established a Fefferman-Graham type expansion in the present setting, and smoothness of the rescaled metric at the conformal boundary [HV24]. We refer in this context also to the gluing and scattering constructions [Hin21, Hin24, GVK17].
@@
A significant aspect of our theorem is that the topology of the spatial slices $\Sigma_{s}$ is $\mathbb{R}\times\mathbb{S}^{2}$ (and thus in the conformal diagram of Figure 2 the hypersurfaces $\Sigma_{s}$ extend to the end points $\iota_{1}^{+}$ and $\iota_{2}^{+}$, see Figure 3). Indeed for compact subsets $K\subset\Sigma_{s}$, with $s\geq s_{0}$ taken sufficiently large, a range of results in the literature imply that the domain of dependence $\mathcal{D}^{+}(K)$ is contained in a perturbation of de Sitter space: [^14] For a longer discussion of the spatially compact setting and its relation to earlier work on de Sitter see [Sch22, Section 1.6]. The de Sitter solution $(\mathbb{H},h)$ can be realized as a hyperboloid $\mathbb{H}$ in $\mathbb{R}^{4+1}$, with metric $h=m\rvert_{\mathbb{H}}$ induced by the ambient Minkowski metric $m$. See also [Vas10] or [Sch21] where the embedding $\mathbb{H}\subset\mathbb{R}^{4+1}$ is used, and its geometric properties are further discussed. In particular the conformal method applies [Fri86, MVK23], and more generally, Ringström proved geodesic completeness from spatially bounded data irrespective of the topology [Rin08].
@@
The gauge (1.13) is crucial for the global existence proof in this setting and leads to a foliation that correctly identifies $\mathcal{I}^{+}$. (In particular, no refoliation is necessary in the continuation argument for Theorem 1.2.) A related problem occurs in the study of stable spacelike singularities, where related ADM gauges have been used to synchronize the singularity [RS18a, FRS23, FL23]; see also [AF25]. In the setting of expanding cosmologies, related gauges have been used in [LvEUW04, Fou22, FMO24].
@@
Since [HV18] provides the stability of the regions $\mathcal{S}_{i}$ on a domain that extends beyond both the cosmological horizon and the event horizon, the result of Hintz-Vasy can in principle be combined with a theorem of Dafermos-Luk on the $C^{0}$-stability of the Cauchy horizon [DL17, Section 1.6]. [^15] For results in spherical symmetry with $\Lambda>0$ in this region see [CGaNS15a, CGaNS15b, CGaNS17]. Further to Corollary 1.3, the combination of all three theorems shows in particular that the Penrose diagram in Fig. 1 is dynamically stable.